OpenStudy (kittiwitti1):

\[2^{\frac{5}{2}} - 2^{\frac{3}{2}}\]can anyone help solve this? like provide an example of the process? answer not needed, just need a start on this the fractions are exponents btw The correct answer given was \[2^{\frac{3}{2}}\] \[\text{Solved, thank you very much!}\] @Abhisar

OpenStudy (abhisar):

I will do it like this \(\sf \huge 2^{\frac{5}{2}} = \frac{1}{2^5}\)

OpenStudy (kittiwitti1):

wut that's possible

OpenStudy (kittiwitti1):

*Is very rusty in algebra.*

OpenStudy (abhisar):

oops...one min

OpenStudy (kittiwitti1):

xD

OpenStudy (abhisar):

I will do it like this \(\sf \huge 2^{\frac{5}{2}} = \sqrt{2}^5\)

Nnesha (nnesha):

;) :)

OpenStudy (kittiwitti1):

o-o

OpenStudy (kittiwitti1):

Yes... The 5 is outside the radical?

Nnesha (nnesha):

\[\huge\sqrt[5]{2}\] ??

OpenStudy (kittiwitti1):

It is? I did it the way Abhi did it. ._.

OpenStudy (abhisar):

That's outside the radical,, it's power of root 5

Nnesha (nnesha):

gotcho

OpenStudy (abhisar):

I will do it like this \(\sf \huge 2^{\frac{5}{2}} = (\sqrt{2})^5=?\)

OpenStudy (kittiwitti1):

Alright then o-o

OpenStudy (kittiwitti1):

\[(\sqrt{2})^{5}-(\sqrt{2})^{3} = (\sqrt{2})^{3}\left((\sqrt{2})^{2}-(\sqrt{2})^{0}\right)\]\[(\sqrt{2})^{3}\left(2-1\right) = (\sqrt{2})^{3}(1) = (\sqrt{2})^{3}\]\[\text{This simplifies to } 2^{\frac{3}{2}}\]\[\text{I have the correct answer. Thank you}\] @Abhisar

OpenStudy (abhisar):

Welcome ^_^

OpenStudy (kittiwitti1):

^-^